OpenStudy (anonymous):

What is the equation of a line passing through (3, 4) and having a slope of -5

OpenStudy (mertsj):

You are given a POINT and the SLOPE so it would be sensible to use POINT-SLOPE form of a linear equation which is: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

now just place the numbers in the equation MertsJ gave you... y-4=-5(x-3) and there you have it :D